OpenStudy (anonymous):

How do I figure out the promblem of 18 - 2q=4?

OpenStudy (anonymous):

4-18/-2

jimthompson5910 (jim_thompson5910):

18-2q = 4 -2q = 4-18 -2q = -14 q = (-14)/(-2) q = 7